Which will have stronger force of attraction-12 protons with 12 electrons or 12 protons with 5 electrons?

Grade:12th pass

2 Answers

Ishan Tarunesh
10 years ago
the force between two charged particle is given by F = k(q1)(q2)/distance squared So we get F is directly proportional to the product of charges considering only the magnitude. If the product is negative then they will attract and if the product is positive they will repel. Now coming to your question the force of attraction between 12 protons and 12 electrons is F = -k (144)/rsquared and the force between 12 proton and 5 electron will be F= -k(60)/r squared. Considering the distances to be equal the force of attraction between 12 proton and 12 electron is more.
Kenny Toijam
10 years ago
Thank you.